A population of N=15 scores has a mean of u=8  One score in the population is changed from X=20 to X=5. What is the value for the new population mean?


1
Expert's answer
2020-07-06T20:22:15-0400

"\mu=\frac{{\sum x_i}} {n}"

"=\frac {(15*8)-20+5}{15}"

=7